How many molecules are present in 2 moles of octane


1
Expert's answer
2021-09-02T02:01:54-0400

A mole of octane contains 6.02×1023 molecules.

Two moles of octane contains "2 \times 6.02 \times 10^{23} = 1.204 \times 10^{24}" molecules.